OpenStudy (milo123):

Help! D? The leg of a right triangle is 4 units and the hypotenuse is 9 units. What is the length, in units, of the other leg of the triangle? square root 97 units 97 units square root 65 units 65 units

OpenStudy (milo123):

9^2=4^2+b^2 81-16=65 @mathmale

OpenStudy (milo123):

@Directrix @mathmale @skullpatrol @Elsa213 check please!

OpenStudy (ashauntay):

its c

OpenStudy (radar):